# Extended-Krylov-subspace methods for trust-region and norm-regularization subproblems

### v2

@@ Appendix C Detailed numerical experiments
name $n$ $\Delta$ $f(x_{*})$ $k_{\mbox{\tiny stop}}$ time facts time prods time
ARWHEAD 5000 1.0 -9.99800000E+03 0 0.00 2 0.00 2 0.00
ARWHEAD 5000 0.5 -9.99800000E+03 1 0.00 3 0.00 1 0.00
BA-L16LS 66462 1.0 -4.00647969E+09 10 1.18 6 1.94 69 2.76
BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 1.43 34 1.42
BA-L16LS 66462 1.0 -4.00647969E+09 10 0.82 6 0.94 69 2.55
BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 0.68 34 1.34
ARWHEAD 5000 1.0 -9.99800000E+03 -1 0.00 2 0.00 2 0.00
ARWHEAD 5000 0.5 -9.99800000E+03 1 0.00 3 0.00 1 0.00
BA-L16LS 66462 1.0 -4.00647969E+09 10 0.84 6 0.94 69 2.60
BA-L16LS 66462 0.5 -1.27074021E+09 10 0.00 10 0.73 34 1.32
BA-L21LS 34134 1.0 -1.43350953E+08 100† 0.88 8 0.39 1129 18.09
BA-L21LS 34134 0.5 -1.27987567E+08 100† 0.00 12 0.27 564 8.95
BA-L49LS 23769 1.0 -2.32581840E+10 6 0.25 8 0.45 27 0.40
BA-L49LS 23769 0.5 -5.82477997E+09 6 0.00 11 0.20 13 0.21
BA-L52LS 192627 1.0 -1.78720629E+10 31 4.65 11 9.18 173 27.82
BA-L52LS 192627 0.5 -1.26606809E+10 31 0.01 16 4.79 86 13.82
BA-L73LS 33753 1.0 -9.88938047E+10 4 0.37 10 0.93 25 0.56
BA-L73LS 33753 0.5 -2.47259422E+10 4 0.00 13 0.35 12 0.29
BDQRTIC 5000 1.0 -4.70328224E+05 4 0.01 4 0.01 13 0.00
BDQRTIC 5000 0.5 -4.37969301E+05 4 0.00 8 0.01 6 0.00
BOX 10000 1.0 -4.99665149E+01 1 0.01 3 0.01 3 0.00
BOX 10000 0.5 -2.49841935E+01 1 0.00 6 0.01 1 0.00
BOXPOWER 20000 1.0 -1.27930301E+05 3 0.01 3 0.02 5 0.00
BOXPOWER 20000 0.5 -7.39842920E+04 3 0.00 6 0.02 2 0.00
BROYDN3DLS 5000 1.0 -5.47141790E+02 6 0.01 3 0.02 19 0.00
BROYDN3DLS 5000 0.5 -2.78620552E+02 6 0.00 6 0.01 9 0.00
BROYDN7D 5000 1.0 -1.04198440E+03 4 0.01 3 0.01 11 0.00
BROYDN7D 5000 0.5 -5.29359803E+02 4 0.00 6 0.01 5 0.00
BROYDNBDLS 5000 1.0 -7.54822794E+03 5 0.02 3 0.01 15 0.01
BROYDNBDLS 5000 0.5 -3.83192245E+03 5 0.00 6 0.01 7 0.00
BRYBND 5000 1.0 -7.54822794E+03 5 0.02 3 0.01 15 0.01
BRYBND 5000 0.5 -3.83192245E+03 5 0.00 6 0.01 7 0.00
CHAINWOO 4000 1.0 -4.18808861E+05 4 0.01 2 0.01 9 0.00
CHAINWOO 4000 0.5 -2.10745351E+05 4 0.00 4 0.01 4 0.00
COSINE 10000 1.0 -7.33802606E+01 5 0.03 2 0.02 11 0.00
COSINE 10000 0.5 -3.63234227E+01 5 0.00 4 0.01 5 0.00
CRAGGLVY 5000 1.0 -2.70357001E+05 2 0.02 3 0.01 5 0.00
CRAGGLVY 5000 0.5 -1.38612789E+05 2 0.00 6 0.01 2 0.00
CURLY10 10000 1.0 -2.55466912E+03 52 0.11 3 0.05 157 0.04
CURLY10 10000 0.5 -6.72371854E+02 52 0.00 6 0.04 78 0.02
CURLY20 10000 1.0 -9.12107858E+03 49 0.11 3 0.05 205 0.07
CURLY20 10000 0.5 -2.35575388E+03 49 0.00 6 0.04 102 0.03
CURLY30 10000 1.0 -1.97300910E+04 41 0.15 3 0.10 235 0.11
CURLY30 10000 0.5 -5.06068395E+03 41 0.00 6 0.06 117 0.04
DIXMAANB 3000 1.0 -1.94571746E+03 4 0.01 3 0.01 11 0.00
DIXMAANB 3000 0.5 -9.82394549E+02 4 0.00 6 0.01 5 0.00
DIXMAANC 3000 1.0 -3.67420974E+03 4 0.02 3 0.01 11 0.00
DIXMAANC 3000 0.5 -1.85594245E+03 4 0.00 6 0.01 5 0.00
DIXMAAND 3000 1.0 -7.40783941E+03 4 0.01 3 0.01 11 0.00
DIXMAAND 3000 0.5 -3.74285037E+03 4 0.00 6 0.01 5 0.00
DIXMAANE1 3000 1.0 -1.03848735E+03 4 0.01 3 0.00 11 0.00
DIXMAANE1 3000 0.5 -5.25114487E+02 4 0.00 6 0.00 5 0.00
DIXMAANF 3000 1.0 -1.83737329E+03 4 0.01 3 0.01 11 0.00
DIXMAANF 3000 0.5 -9.28137690E+02 4 0.00 6 0.01 5 0.00
DIXMAANG 3000 1.0 -3.56193345E+03 4 0.01 3 0.01 11 0.00
DIXMAANG 3000 0.5 -1.79971801E+03 4 0.00 6 0.01 5 0.00
DIXMAANH 3000 1.0 -7.28769893E+03 4 0.01 3 0.01 11 0.00
DIXMAANH 3000 0.5 -3.68269064E+03 4 0.00 6 0.01 5 0.00
DIXMAANI1 3000 1.0 -1.00064454E+03 4 0.01 3 0.00 11 0.00
DIXMAANI1 3000 0.5 -5.06141189E+02 4 0.00 6 0.00 5 0.00
DIXMAANJ 3000 1.0 -1.79984433E+03 4 0.01 3 0.01 11 0.00
DIXMAANJ 3000 0.5 -9.09324798E+02 4 0.00 6 0.01 5 0.00
DIXMAANK 3000 1.0 -3.52376478E+03 4 0.01 3 0.01 11 0.00
DIXMAANK 3000 0.5 -1.78058448E+03 4 0.00 6 0.01 5 0.00
DIXMAANL 3000 1.0 -7.24829529E+03 4 0.01 3 0.01 11 0.00
DIXMAANL 3000 0.5 -3.66293885E+03 4 0.00 6 0.01 5 0.00
DIXMAANM1 3000 1.0 -4.27493497E+02 5 0.01 3 0.01 13 0.00
DIXMAANM1 3000 0.5 -2.16348948E+02 5 0.00 6 0.01 6 0.00
DIXMAANN 3000 1.0 -9.97842749E+02 5 0.01 3 0.01 13 0.00
DIXMAANN 3000 0.5 -5.05235340E+02 5 0.00 6 0.01 6 0.00
DIXMAANO 3000 1.0 -1.90145642E+03 5 0.01 3 0.01 13 0.00
DIXMAANO 3000 0.5 -9.63139983E+02 5 0.00 6 0.01 6 0.00
DIXMAANP 3000 1.0 -3.85352076E+03 5 0.01 3 0.01 13 0.00
DIXMAANP 3000 0.5 -1.95233964E+03 5 0.00 6 0.01 6 0.00
DIXON3DQ 10000 1.0 -4.35180402E+00 19 0.02 5 0.03 45 0.01
DIXON3DQ 10000 0.5 -2.48148299E+00 19 0.00 8 0.02 22 0.00
DQDRTIC 5000 1.0 -8.50546818E+04 3 0.00 2 0.00 7 0.00
DQDRTIC 5000 0.5 -4.25775883E+04 3 0.00 4 0.00 3 0.00
DQRTIC 5000 1.0 -1.33489191E+13 2 0.00 2 0.00 5 0.00
DQRTIC 5000 0.5 -6.67448869E+12 2 0.00 4 0.00 2 0.00
EDENSCH 2000 1.0 -9.90061935E+04 3 0.00 2 0.00 7 0.00
EDENSCH 2000 0.5 -4.96303271E+04 3 0.00 4 0.00 3 0.00
EG2 1000 1.0 -1.73066127E+02 0 0.00 2 0.00 1 0.00
EG2 1000 0.5 -1.64667237E+02 1 0.00 4 0.00 0 0.00
EIGENALS 2550 1.0 -1.01221801E+03 7 0.75 3 0.85 25 0.04
EIGENALS 2550 0.5 -4.77552997E+02 7 0.00 6 0.79 12 0.02
EIGENBLS 2550 1.0 -3.52937908E+01 8 0.70 3 0.89 25 0.04
EIGENBLS 2550 0.5 -1.80972156E+01 8 0.00 5 0.62 12 0.02
EIGENCLS 2652 1.0 -4.70591220E+02 9 0.78 3 0.91 31 0.04
EIGENCLS 2652 0.5 -2.33514688E+02 9 0.00 6 0.93 15 0.02
ENGVAL1 5000 1.0 -8.67081566E+03 3 0.01 2 0.01 9 0.00
ENGVAL1 5000 0.5 -4.35940622E+03 3 0.00 4 0.01 4 0.00
EXTROSNB 1000 1.0 -3.66203611E+04 4 0.00 3 0.00 11 0.00
EXTROSNB 1000 0.5 -1.86350898E+04 4 0.00 6 0.00 5 0.00
FLETBV3M 5000 1.0 -4.36166522E+01 3 0.01 2 0.01 7 0.00
FLETBV3M 5000 0.5 -2.18337960E+01 3 0.00 4 0.01 3 0.00
FLETCBV2 5000 1.0 -3.88209445E-06 13 0.01 4 0.01 9737 412.07
FLETCBV2 5000 0.5 -2.02683952E-06 16 0.00 8 0.01 4868 0.53
FLETCBV3 5000 1.0 -4.36172752E+01 3 0.01 2 0.01 7 0.04
FLETCBV3 5000 0.5 -2.18341109E+01 3 0.00 4 0.01 3 0.00
FLETCHBV 5000 1.0 -2.74868337E+09 3 0.01 2 0.01 9 0.00
FLETCHBV 5000 0.5 -1.37709990E+09 3 0.00 4 0.01 4 0.00
FLETCHCR 1000 1.0 -1.08786732E+01 1 0.00 3 0.00 3 0.00
FLETCHCR 1000 0.5 -1.05114880E+01 1 0.00 7 0.00 1 0.00
FMINSRF2 5625 1.0 -2.74210232E-01 44 0.04 4 0.03 107 0.03
FMINSRF2 5625 0.5 -1.42608286E-01 44 0.00 8 0.03 53 0.01
FMINSURF 5625 1.0 -2.74396123E-01 44 5.28 4 8.28 107 0.03
FMINSURF 5625 0.5 -1.42697922E-01 44 0.00 8 8.05 53 0.01
FREUROTH 5000 1.0 -5.51793805E+04 3 0.01 3 0.01 7 0.00
FREUROTH 5000 0.5 -2.75854293E+04 3 0.00 6 0.01 3 0.00
GENHUMPS 5000 1.0 -6.64118303E+03 6 0.03 3 0.02 13 0.00
GENHUMPS 5000 0.5 -3.16552923E+03 6 0.00 6 0.01 6 0.00
INDEF 5000 1.0 -2.10493952E+03 6 0.01 103‡ 0.12 13 0.00
INDEF 5000 0.5 -5.26825583E+02 6 0.00 207‡ 0.25 6 0.00
INDEFM 100000 1.0 -4.20748078E+04 3 0.09 11 0.38 11 0.05
INDEFM 100000 0.5 -1.05192638E+04 3 0.00 16 0.23 5 0.03
JIMACK 3549 1.0 -4.37368349E-01 100† 0.45 26 0.57 3201 17.07
JIMACK 3549 0.5 -2.41925217E-01 100† 0.00 31 0.14 1600 0.62
LIARWHD 5000 1.0 -4.61798034E+05 1 0.00 3 0.00 3 0.01
LIARWHD 5000 0.5 -2.36034628E+05 1 0.00 6 0.00 1 0.00
MODBEALE 20000 1.0 -3.04981491E+05 2 0.05 2 0.03 7 0.01
MODBEALE 20000 0.5 -1.52960464E+05 2 0.00 4 0.03 3 0.00
MOREBV 5000 1.0 -7.72418940E-14 9 0.01 107‡ 0.13 295 0.17
MOREBV 5000 0.5 -7.72418940E-14 54 0.03 116 0.02 294 0.02
MSQRTALS 1024 1.0 -3.70040042E+02 4 0.16 4 0.14 27 0.01
[…]

### v3

@@ Appendix D Detailed numerical experiments
(a) (b) (a) (b) (a) (b)
name $n$ $\Delta$ $f(x_{*})$ $k_{\mbox{\tiny s}}$ time $k_{\mbox{\tiny s}}$ time fac time fac time pr time pr time
ARWHEAD 5000 10 -9.99800000E+03 0 0.01 2 0.00 2 0.00
ARWHEAD 5000 .1 -3.59936000E+03 1 0.00 0 0.01 4 0.00 2 0.00 1 0.00 2 0.00
ARWHEAD 5000 .01 -3.95930600E+02 1 0.00 1 0.00 6 0.00 4 0.00 1 0.00 1 0.00
BA-L16LS 66462 10 -3.46265942E+11 12 0.91 9 1.28 75 2.11
BA-L16LS 66462 1 -4.00647969E+09 12 0.00 10 0.82 13 0.79 6 1.06 37 1.08 69 1.96
BA-L16LS 66462 .1 -2.71798703E+08 12 0.00 10 0.00 19 0.98 12 1.02 37 1.09 34 1.01
BA-L21LS 34134 10 -1.09802044E+09 284 11.60 15 0.70 1999 22.79
BA-L21LS 34134 1 -1.43350953E+08 284 0.15 196 3.34 21 0.39 8 0.43 999 11.38 1129 13.46
BA-L21LS 34134 1 -1.43350953E+08 284 0.15 196 0.05 22 0.15 14 0.39 999 11.36 564 6.62
BA-L49LS 23769 10 -2.32220041E+12 6 0.25 9 0.57 29 0.32
BA-L49LS 23769 1 -2.32581840E+10 6 0.00 6 0.24 12 0.25 8 0.59 14 0.17 27 0.31
BA-L49LS 23769 .1 -2.36543531E+08 6 0.00 6 0.00 15 0.26 11 0.28 14 0.17 13 0.17
BA-L52LS 192627 10 -5.93753593E+11 34 5.24 9 8.15 221 25.43
BA-L52LS 192627 1 -1.78720629E+10 34 0.01 31 4.98 14 5.33 11 10.32 110 12.86 173 20.84
BA-L52LS 192627 .1 -5.16366984E+09 34 0.01 31 0.01 18 4.42 15 4.81 110 12.84 86 10.49
BA-L73LS 33753 10 -9.88855550E+12 5 0.34 12 1.12 27 0.44
BA-L73LS 33753 1 -9.88938047E+10 5 0.00 4 0.32 15 0.42 10 0.97 13 0.24 25 0.41
BA-L73LS 33753 .1 -9.90136989E+08 5 0.00 4 0.00 18 0.42 13 0.46 13 0.24 12 0.22
BDQRTIC 5000 10 -6.53953444E+05 6 0.00 3 0.01 29 0.01
BDQRTIC 5000 1 -4.70328224E+05 6 0.00 4 0.00 7 0.02 4 0.02 14 0.00 13 0.00
BDQRTIC 5000 .1 -1.37454488E+05 6 0.00 4 0.00 10 0.01 7 0.01 14 0.00 6 0.00
BOX 10000 10 -4.99340865E+02 1 0.03 3 0.01 3 0.00
BOX 10000 1 -4.99665149E+01 1 0.00 1 0.04 6 0.01 3 0.02 1 0.00 3 0.00
BOX 10000 .1 -4.99704061E+00 1 0.00 1 0.00 9 0.01 6 0.01 1 0.00 1 0.00
BOXPOWER 20000 10 -1.76398331E+05 24 0.15 6 0.04 9 0.00
BOXPOWER 20000 1 -1.27930301E+05 24 0.00 2 0.18 9 0.02 3 0.02 4 0.00 5 0.00
BOXPOWER 20000 .1 -1.63999277E+04 24 0.00 2 0.00 11 0.02 5 0.02 4 0.00 2 0.00
BROYDN3DLS 5000 10 -3.66408186E+03 6 0.00 3 0.01 45 0.00
BROYDN3DLS 5000 1 -5.47141790E+02 6 0.00 6 0.01 6 0.01 3 0.01 22 0.00 19 0.00
BROYDN3DLS 5000 .1 -5.65333513E+01 6 0.00 6 0.00 9 0.01 6 0.01 22 0.00 9 0.00
BROYDN7D 5000 10 -7.40764118E+03 7 0.00 3 0.01 31 0.00
BROYDN7D 5000 1 -1.04198440E+03 7 0.00 4 0.01 6 0.01 3 0.01 15 0.00 11 0.00
BROYDN7D 5000 .1 -1.07210780E+02 7 0.00 4 0.00 8 0.01 5 0.01 15 0.00 5 0.00
BROYDNBDLS 5000 10 -5.46819825E+04 4 0.01 4 0.02 31 0.01
BROYDNBDLS 5000 1 -7.54822794E+03 5 0.00 5 0.01 7 0.02 3 0.02 15 0.00 15 0.00
BROYDNBDLS 5000 .1 -7.75634313E+02 5 0.00 5 0.00 9 0.01 5 0.01 15 0.00 7 0.00
BRYBND 5000 10 -5.46819825E+04 4 0.01 4 0.02 31 0.01
BRYBND 5000 1 -7.54822794E+03 5 0.00 5 0.01 7 0.02 3 0.02 15 0.00 15 0.00
BRYBND 5000 .1 -7.75634313E+02 5 0.00 5 0.00 9 0.01 5 0.01 15 0.00 7 0.00
CHAINWOO 4000 10 -3.70797995E+06 5 0.00 3 0.01 15 0.00
CHAINWOO 4000 1 -4.18808861E+05 5 0.00 4 0.00 5 0.01 2 0.00 7 0.00 9 0.00
CHAINWOO 4000 .1 -4.23637069E+04 5 0.00 4 0.00 7 0.01 4 0.00 7 0.00 4 0.00
COSINE 10000 10 -8.65819784E+02 12 0.01 3 0.01 27 0.00
COSINE 10000 1 -7.33802606E+01 12 0.00 5 0.01 5 0.01 2 0.01 13 0.00 11 0.00
COSINE 10000 .1 -7.20601140E+00 12 0.00 5 0.00 7 0.01 4 0.01 13 0.00 5 0.00
CRAGGLVY 5000 10 -1.47282263E+06 2 0.00 6 0.01 11 0.00
CRAGGLVY 5000 1 -2.70357001E+05 2 0.00 2 0.00 9 0.01 3 0.01 5 0.00 5 0.00
CRAGGLVY 5000 .1 -2.82720573E+04 2 0.00 2 0.00 11 0.01 5 0.01 5 0.00 2 0.00
CURLY10 10000 10 -2.43341009E+05 62 0.09 3 0.08 519 0.10
CURLY10 10000 1 -2.55466912E+03 62 0.00 52 0.07 7 0.05 3 0.05 259 0.05 157 0.03
CURLY10 10000 .1 -3.76837984E+01 62 0.00 52 0.00 10 0.05 6 0.05 259 0.05 78 0.01
CURLY20 10000 10 -8.84977907E+05 36 0.07 4 0.13 677 0.21
CURLY20 10000 1 -9.12107858E+03 49 0.02 49 0.09 8 0.11 3 0.10 338 0.09 205 0.05
CURLY20 10000 .1 -1.18405527E+02 49 0.00 49 0.00 11 0.10 6 0.09 338 0.09 102 0.03
CURLY30 10000 10 -1.92700208E+06 26 0.08 4 0.18 769 0.28
CURLY30 10000 1 -1.97300910E+04 40 0.02 40 0.10 8 0.14 3 0.14 384 0.13 235 0.09
CURLY30 10000 .1 -2.43499809E+02 40 0.00 40 0.00 11 0.12 6 0.14 384 0.14 117 0.04
DIXMAANB 3000 10 -1.60339163E+04 7 0.00 3 0.01 23 0.00
DIXMAANB 3000 1 -1.94571746E+03 7 0.00 4 0.00 6 0.01 3 0.01 11 0.00 11 0.00
DIXMAANB 3000 .1 -1.98005001E+02 7 0.00 4 0.00 8 0.01 5 0.01 11 0.00 5 0.00
DIXMAANC 3000 10 -2.99803804E+04 7 0.00 3 0.01 25 0.00
DIXMAANC 3000 1 -3.67420974E+03 7 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
DIXMAANC 3000 .1 -3.74203237E+02 7 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
DIXMAAND 3000 10 -6.01051338E+04 8 0.00 3 0.01 25 0.00
DIXMAAND 3000 1 -7.40783941E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
DIXMAAND 3000 .1 -7.54800526E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
DIXMAANE1 3000 10 -8.27329007E+03 6 0.00 3 0.00 21 0.00
DIXMAANE1 3000 1 -1.03848735E+03 6 0.00 4 0.00 6 0.00 3 0.00 10 0.00 11 0.00
DIXMAANE1 3000 .1 -1.05962269E+02 6 0.00 4 0.00 8 0.00 5 0.00 10 0.00 5 0.00
DIXMAANF 3000 10 -1.49809432E+04 8 0.00 3 0.01 25 0.00
DIXMAANF 3000 1 -1.83737329E+03 8 0.00 4 0.01 6 0.01 3 0.01 12 0.00 11 0.00
DIXMAANF 3000 .1 -1.87140059E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
DIXMAANG 3000 10 -2.88885934E+04 8 0.00 3 0.01 25 0.00
DIXMAANG 3000 1 -3.56193345E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
DIXMAANG 3000 .1 -3.62944535E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
DIXMAANH 3000 10 -5.89358136E+04 8 0.00 3 0.01 25 0.00
DIXMAANH 3000 1 -7.28769893E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
DIXMAANH 3000 .1 -7.42754250E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
DIXMAANI1 3000 10 -7.91454338E+03 6 0.00 3 0.00 21 0.00
DIXMAANI1 3000 1 -1.00064454E+03 6 0.00 4 0.00 6 0.00 3 0.00 10 0.00 11 0.00
DIXMAANI1 3000 .1 -1.02159327E+02 6 0.00 4 0.00 8 0.00 5 0.00 10 0.00 5 0.00
DIXMAANJ 3000 10 -1.46232627E+04 8 0.00 3 0.01 25 0.00
DIXMAANJ 3000 1 -1.79984433E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
DIXMAANJ 3000 .1 -1.83369741E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
DIXMAANK 3000 10 -2.85248120E+04 8 0.01 3 0.01 25 0.00
DIXMAANK 3000 1 -3.52376478E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
DIXMAANK 3000 .1 -3.59109964E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
DIXMAANL 3000 10 -5.85599682E+04 8 0.00 3 0.01 25 0.00
DIXMAANL 3000 1 -7.24829529E+03 8 0.00 4 0.00 6 0.01 3 0.01 12 0.00 11 0.00
DIXMAANL 3000 .1 -7.38795901E+02 8 0.00 4 0.00 8 0.01 5 0.01 12 0.00 5 0.00
DIXMAANM1 3000 10 -3.35446625E+03 9 0.00 3 0.00 33 0.00
DIXMAANM1 3000 1 -4.27493497E+02 9 0.00 5 0.00 6 0.00 3 0.00 16 0.00 13 0.00
DIXMAANM1 3000 .1 -4.36867895E+01 9 0.00 5 0.00 8 0.00 5 0.00 16 0.00 6 0.00
DIXMAANN 3000 10 -7.76833364E+03 10 0.00 3 0.01 33 0.00
DIXMAANN 3000 1 -9.97842749E+02 10 0.00 5 0.00 6 0.01 3 0.01 16 0.00 13 0.00
DIXMAANN 3000 .1 -1.02059568E+02 10 0.00 5 0.00 8 0.01 5 0.01 16 0.00 6 0.00
DIXMAANO 3000 10 -1.46773625E+04 11 0.00 3 0.01 33 0.00
DIXMAANO 3000 1 -1.90145642E+03 11 0.00 5 0.00 6 0.01 3 0.01 16 0.00 13 0.00
DIXMAANO 3000 .1 -1.94618568E+02 11 0.00 5 0.00 8 0.01 5 0.01 16 0.00 6 0.00
DIXMAANP 3000 10 -2.96054573E+04 11 0.00 3 0.01 35 0.00
DIXMAANP 3000 1 -3.85352076E+03 11 0.00 5 0.00 6 0.01 3 0.01 17 0.00 13 0.00
DIXMAANP 3000 .1 -3.94570550E+02 11 0.00 5 0.00 8 0.01 5 0.01 17 0.00 6 0.00
DIXON3DQ 10000 10 -7.95918012E+00 89 0.14 9 0.03 2359 0.23
DIXON3DQ 10000 1 -4.35180402E+00 89 0.01 19 0.01 15 0.02 5 0.02 1179 0.10 45 0.00
DIXON3DQ 10000 .1 -5.50941460E-01 89 0.01 19 0.00 18 0.01 8 0.01 1179 0.11 22 0.00
DQDRTIC 5000 10 -8.32457765E+05 3 0.00 3 0.00 9 0.00
DQDRTIC 5000 1 -8.50546818E+04 3 0.00 3 0.00 5 0.00 2 0.00 4 0.00 7 0.00
DQDRTIC 5000 .1 -8.52355726E+03 3 0.00 3 0.00 7 0.00 4 0.00 4 0.00 3 0.00
DQRTIC 5000 10 -1.33478697E+14 3 0.00 2 0.00 5 0.00
DQRTIC 5000 1 -1.33489191E+13 3 0.00 2 0.00 4 0.00 2 0.00 2 0.00 5 0.00
DQRTIC 5000 .1 -1.33490240E+12 3 0.00 2 0.00 6 0.00 4 0.00 2 0.00 2 0.00
EDENSCH 2000 10 -9.44259112E+05 5 0.00 3 0.00 13 0.00
EDENSCH 2000 1 -9.90061935E+04 5 0.00 3 0.00 5 0.00 2 0.00 6 0.00 7 0.00
EDENSCH 2000 .1 -9.94642228E+03 5 0.00 3 0.00 7 0.00 4 0.00 6 0.00 3 0.00
EG2 1000 10 -1.73066127E+02 0 0.00 2 0.00 1 0.00
EG2 1000 .1 -4.97676498E+01 1 0.00 0 0.00 4 0.00 2 0.00 0 0.00 1 0.00
EG2 1000 .01 -5.35553453E+00 1 0.00 1 0.00 6 0.00 4 0.00 0 0.00 0 0.00
EIGENALS 2550 10 -2.15497836E+04 8 0.50 4 1.12 75 0.08
EIGENALS 2550 1 -1.01221801E+03 8 0.00 7 0.46 7 0.91 3 1.07 37 0.04 25 0.03
EIGENALS 2550 .1 -9.10252017E+01 8 0.00 7 0.00 10 0.90 6 1.14 37 0.04 12 0.02
[…]
EIGENBLS 2550 1 -3.52937908E+01 49 0.00 8 0.55 116 0.91 3 1.07 301 0.34 25 0.03
EIGENBLS 2550 .1 -3.70678144E+00 49 0.00 8 0.00 119 0.90 6 0.96 301 0.32 12 0.02
EIGENCLS 2652 10 -9.33034837E+03 9 0.57 4 1.26 83 0.10
EIGENCLS 2652 1 -4.70591220E+02 9 0.00 9 0.54 7 0.99 3 1.08 41 0.05 31 0.04
EIGENCLS 2652 .1 -4.70009302E+01 9 0.00 9 0.00 9 0.77 5 0.81 41 0.05 15 0.02
ENGVAL1 5000 10 -7.80687659E+04 5 0.00 3 0.01 15 0.00
ENGVAL1 5000 1 -8.67081566E+03 5 0.00 3 0.00 5 0.01 2 0.01 7 0.00 9 0.00
ENGVAL1 5000 .1 -8.75720987E+02 5 0.00 3 0.00 7 0.01 4 0.01 7 0.00 4 0.00
EXTROSNB 1000 10 -2.49243249E+05 6 0.00 3 0.00 27 0.00
EXTROSNB 1000 1 -3.66203611E+04 6 0.00 4 0.00 6 0.00 3 0.00 13 0.00 11 0.00
EXTROSNB 1000 .1 -3.77900359E+03 6 0.00 4 0.00 8 0.00 5 0.00 13 0.00 5 0.00
FLETBV3M 5000 10 -4.27009603E+02 4 0.00 3 0.01 11 0.00
FLETBV3M 5000 1 -4.36166522E+01 4 0.00 3 0.00 5 0.01 2 0.01 5 0.00 7 0.00
FLETBV3M 5000 .1 -4.37083490E+00 4 0.00 3 0.00 7 0.01 4 0.01 5 0.00 3 0.00
FLETCBV2 5000 10 -1.81067238E-05 0 0.00 2 0.01 4869 0.47
FLETCBV2 5000 1 -3.88209445E-06 8 0.00 8 0.01 6 0.01 4 0.01 4868 0.44 9737 1.47
FLETCBV2 5000 .1 -4.25968981E-07 16 0.00 16 0.00 10 0.01 8 0.01 4868 0.45 4868 0.47
FLETCBV3 5000 10 -4.27014602E+02 4 0.00 3 0.01 11 0.00
FLETCBV3 5000 1 -4.36172752E+01 4 0.00 3 0.00 5 0.01 2 0.01 5 0.00 7 0.00
FLETCBV3 5000 .1 -4.37089841E+00 4 0.00 3 0.00 7 0.01 4 0.01 5 0.00 3 0.00
FLETCHBV 5000 10 -2.64954410E+10 4 0.00 3 0.01 13 0.00
FLETCHBV 5000 1 -2.74868337E+09 4 0.00 3 0.00 5 0.01 2 0.01 6 0.00 9 0.00
FLETCHBV 5000 .1 -2.75861358E+08 4 0.00 3 0.00 7 0.01 4 0.01 6 0.00 4 0.00
FLETCHCR 1000 10 -1.08811881E+01 0 0.00 2 0.00 2 0.00
FLETCHCR 1000 1 -1.08786732E+01 1 0.00 1 0.00 5 0.00 3 0.00 1 0.00 3 0.00
FLETCHCR 1000 .1 -5.31285550E+00 1 0.00 1 0.00 8 0.00 6 0.00 1 0.00 1 0.00
FMINSRF2 5625 10 -2.48385848E+00 130 0.37 4 0.04 331 0.07
FMINSRF2 5625 1 -2.74210232E-01 130 0.01 42 0.03 8 0.03 4 0.04 165 0.04 107 0.02
FMINSRF2 5625 1 -2.74210232E-01 130 0.01 42 0.00 9 0.02 8 0.03 165 0.04 53 0.01
FMINSURF 5625 10 -2.48582974E+00 129 9.32 4 8.90 331 0.09
FMINSURF 5625 1 -2.74396123E-01 129 0.01 42 4.93 8 8.57 4 9.22 165 0.04 107 0.02
FMINSURF 5625 1 -2.74396123E-01 129 0.01 42 0.00 9 3.25 8 9.10 165 0.04 53 0.01
FREUROTH 5000 10 -5.53358711E+05 4 0.00 3 0.01 11 0.00
FREUROTH 5000 1 -5.51793805E+04 4 0.00 3 0.00 6 0.01 3 0.01 5 0.00 7 0.00
FREUROTH 5000 .1 -5.51640621E+03 4 0.00 3 0.00 8 0.01 5 0.01 5 0.00 3 0.00
GENHUMPS 5000 10 -1.22237034E+05 13 0.01 3 0.01 21 0.01
GENHUMPS 5000 1 -6.64118303E+03 13 0.00 6 0.00 6 0.01 3 0.01 10 0.00 13 0.00
GENHUMPS 5000 .1 -6.08296147E+02 13 0.00 6 0.00 8 0.01 5 0.01 10 0.00 6 0.00
[…]
INDEFM 100000 1 -4.20748078E+04 184 1.57 70 0.38 120 0.35 11 0.37 5 0.01 11 0.03
INDEFM 100000 1 -4.20748078E+04 184 0.09 70 0.01 121 0.08 15 0.18 5 0.01 5 0.01
[…]
JIMACK 3549 1 -4.37368349E-01 24 0.05 24 0.05 33 0.04 34 0.15 1723 0.85 1600 0.96
LIARWHD 5000 10 -2.76920956E+06 1 0.01 4 0.01 3 0.00
LIARWHD 5000 1 -4.61798034E+05 1 0.00 1 0.01 7 0.00 3 0.00 1 0.00 3 0.00
LIARWHD 5000 .1 -4.80286236E+04 1 0.00 1 0.00 9 0.00 5 0.00 1 0.00 1 0.00
MODBEALE 20000 10 -2.88071792E+06 3 0.01 3 0.03 9 0.01
MODBEALE 20000 1 -3.04981491E+05 3 0.00 2 0.01 5 0.02 2 0.03 4 0.00 7 0.02
MODBEALE 20000 .1 -3.06672478E+04 3 0.00 2 0.00 7 0.02 4 0.03 4 0.00 3 0.01
MOREBV 5000 10 -7.72418940E-14 0 0.00 2 0.01 295 0.02
[…]
MOREBV 5000 .1 -7.72418940E-14 8 0.00 8 0.00 119 0.02 117 0.02 294 0.02 294 0.02
MSQRTALS 1024 10 -7.41223812E+03 16 0.14 4 0.15 85 0.03
MSQRTALS 1024 1 -3.70040042E+02 16 0.00 4 0.12 8 0.14 4 0.15 42 0.02 27 0.01
MSQRTALS 1024 .1 -3.36456076E+01 16 0.00 4 0.00 11 0.12 7 0.13 42 0.02 13 0.00
MSQRTBLS 1024 10 -7.39900251E+03 16 0.14 4 0.15 85 0.03
MSQRTBLS 1024 1 -3.69312117E+02 16 0.00 4 0.12 8 0.15 4 0.16 42 0.02 27 0.01
MSQRTBLS 1024 .1 -3.35861466E+01 16 0.00 4 0.00 11 0.12 7 0.13 42 0.02 13 0.01
NCB20 5010 10 -2.55084157E+03 13 0.02 4 0.04 57 0.09
[…]